A stuntman weighing 70kg is launched into the air with a velocity of 35 m/s.
a) What is his height and velocity after being in the air 2 seconds?
b) What is his height and velocity after being in the air 5 seconds?
c) What is the highest point he can reach?
d) When will he land on the air mattress after his flight?
Answer #1

. U=35m/s a=-9.8 m/ a v=utat » 35-9.8x2 = 15-4 m/ szre > (15:47-3571 50.4m. - 229-8 Voltat = 35-9.8x5=-14m/s. sa vad cea - fl